Assume that General Dynamics, which manufactures high-technology instruments for spacecraft, is considering the sale of a navigational unit to a government agency in India that wishes to launch its own communications satellite. The government agency plans to purchase 8 units, although it would also consider buying 16 units. General Dynamics has started a chart relating labor time required to units produced.

Unit Produced (X) Time Required to Produce
the Xth Unit
1 16,000 hours
2 11,200 hours
4 7,840 hours
8 ?
16 ?

Required:

a. Compute the labor time required to produce 8 and 16 units. (Round your answers to the nearest whole number.)


b. Assume that labor time costs $140 per hour. Compare the cost of producing the 1st unit to the cost of producing the 16th unit. What is the percentage of the cost of the 16th unit to the cost of the 1st unit? (Do not round your intermediate calculations. Enter your answer as a percentage rounded to 2 decimal places (i.e., 32.21))
Answer #1

Answers

  • [a]

Unit Produced (X)

Time Required to Produce

the Xth Unit

1

16,000 hours

2

11,200 hours

[11200/16000 = 70%]

4

7,840 hours

[11200 x 70%]

8

5488 hours

[7840 x 70%]

16

3842 hours

[5488 x 70%]

  • [b]

Unit Produced (X)

Time Required to Produce

Total Labor Cost

the Xth Unit

[Hours x Units x $ 140]

1

16,000 hours

$2,240,000

=16000*1*140

2

11,200 hours

$3,136,000

=11200*2*140

4

7,840 hours

$4,390,400

=7840*4*140

8

5488 hours

$6,146,560

=5488*8*140

16

3842 hours

$8,606,080

=3842*16*140

Unit Produced (X)

Time Required to Produce

Total Labor Cost

Cost of producing the unit

the Xth Unit

[Hours x Units x $ 140]

[Hours x $ 140]

1

16,000 hours

$2,240,000

$2,240,000

2

11,200 hours

$3,136,000

$1,568,000

4

7,840 hours

$4,390,400

$1,097,600

8

5488

$6,146,560

$768,320

16

3842

$8,606,080

$537,880

>Percentage of 16th unit cost to 1st unit
= ($537880 / $2240000) x 100
= 24.01%